# Fluctuations of Young diagrams for symplectic groups and semiclassical orthogonal polynomials

### v2

@@ 2.1. Symplectic polynomials
$\sum\limits_{a=0}^{K}\binom{K}{a}p^{a}(1-p)^{K-a}K_{m}(a,p,K)K_{n}(a,p,K)=\frac{(-1)^{n}n!}{(-K)_{l}}\left(\frac{1-p}{p}\right)^{n}\delta_{mn}=\binom{K}{n}^{-1}\left(\frac{1-p}{p}\right)^{n}\delta_{nm},$
@@
There is a well known representation of the Krawtchouk polynomials as hypergeometric functions [32]
@@
The monic version of the $K_{m}(a,p,K)$ is defined as (9.11.4 in [32])
@@
$P_{m}(a,p,K)=(-1)^{m}m!p^{m}\binom{K}{m}K_{m}(a,p,K)=(-1)^{m}m!p^{m}\binom{K}{m}\,_{2}F_{1}\left(-m,-a,-K;\frac{1}{p}\right)$ (2.3)
@@
Since the weight $\widetilde{W}(u^{2})$, defined by (1.7) differs from the weight of Krawthouk polynomials $W$ with $K=2n+2k$ only by a factor $u^{2}$, monic orthogonal polynomials $G_{m}$ are expressed in terms of the polynomials (2.4) by the formula (Christoffel transformation, see [33], sect. 2.7, Theorem 2.7.1)
@@
where $\Lambda_{m}=||G_{m}||$ can be computed by the recurrence (see next section). The monic symplectic polynomials $\{G_{m}(u)\}$ have the following hypergeometric realization for even $m$
@@
$-\frac{\,{}_{2}F_{1}\left(-m-2,-K/2,-K;\frac{1}{p}\right)}{\,{}_{2}F_{1}\left(-m,-K/2,-K;\frac{1}{p}\right)}\,_{2}F_{1}\left(-m,-K/2-nu,-K;\frac{1}{p}\right)\bigg{)},$ (2.12)
@@
2.2. QR algorithm and three term relation’s coefficients
Using the results from [24], one can derive a three-term recurrence relation for symplectic polynomials. This can be achieved by employing the fact that the Christoffel transformation provides a method for constructing orthogonal polynomials with respect to the measure $x^{2}W(x)dx$ from those that are orthogonal with respect to the positive Borel measure $W(x)dx$.
@@ 2.2. QR algorithm and three term relation’s coefficients
[…]
@@
[…]
@@
[…]
@@
Now we can determine the coefficients of the three term relation for the polynomials $\{\pi_{m}(x)\}$ from three term relation coefficients $\{\tilde{\alpha}_{m},\tilde{\beta}_{m}\}$ of the monic polynomials $\{\widetilde{K}_{m}\}$ (see [32])
$x\widetilde{K}_{m}(x)=\widetilde{K}_{m+1}(x)+\tilde{\alpha}_{m}\widetilde{K}_{m}(x)+\tilde{\beta}_{m}\widetilde{K}_{m-1}(x)=$ (2.18)
$=\widetilde{K}_{m+1}(x)+[p(K-m)+m(1-p)-K/2]\widetilde{K}_{m}(x)+\frac{m(K-m+1)}{n^{2}}p(1-p)\widetilde{K}_{m-1}(x).$
@@
$a_{m}=\frac{\kappa_{m-1}}{n\kappa_{m}}=-\frac{\big{[}m(K-m+1)(p(1-p))\big{]}^{1/2}}{n}=-\sqrt{\tilde{\beta}_{m}},$
$b_{m}=\left[p(K-m)+m(1-p)-K/2\right]=\tilde{\alpha}_{m}.$
@@
Following the Theorem 3.3 (see [24]), consider the Jacobi matrix of the three term relation coefficient for the orthonormal polynomials with the measure $\mu$. Denote by $\omega$ the natural mapping from the set of all normalized, real, positive Borel measures $\mathcal{B}$ to the set of all singly-infinite, symmetric, tridiagonal irreducible matrices $\mathcal{J}$
@@
that takes every member of $\mathcal{B}$ to the corresponding Jacobi matrix. By the Favard theorem [34] this mapping is invertible if and only if the underlying Hamburger problem is determinate [35]. By the inverse mapping we mean $\omega^{-1}:\mathcal{J}\to\mathcal{B}^{*}$, where $\mathcal{B}^{*}$ is the quotient by the equivalence classes of the measures with identical moment sequences. In our case the measure $d\mu(x)$ defined above has the discrete and finite support, which means that the moment problem for $\text{Im}(\omega)$ is always determinate and the equivalence class of the element $\omega^{-1}(J)$ consists of a single element (see [36], [37], [38]). Then the QR algorithm gives the reversible arrows in the diagram
${J}$${\widehat{J}}$${d\mu}$${d\tilde{\mu}}$$\scriptstyle{\tau}$$\scriptstyle{\omega}$$\scriptstyle{\rho}$$\scriptstyle{\omega}$
@@
$QR$ factorization of Jacobi matrix $J$ is its decomposition into the product $J=Q_{1}^{T}Q_{2}^{T}...Q_{n-1}^{T}R$, where $Q_{i}$ are block-diagonal antisymmetric matrices, which turn the subdiagonal elements to zero (Givens rotation). For example, letting $r_{0}=\sqrt{a_{1}^{2}+b_{0}^{2}}$, we take $Q_{1}$ to be identity except that the upper left $2\times 2$ block is
@@
The trailing coefficients $b_{i}^{*}$ and $a_{i+1}^{*}$, obtained by applying $i$ successive Givens rotations to $J$, serve as the parameters for the subsequent rotation $Q_{i+1}$. The nontrivial part of $Q_{i+1}$, in its $(k+1)$st and $(k+2)$nd rows and columns, has the form
@@
Denoting $n^{-2m}\kappa_{m}^{-2}=L_{m}$ in (2.5), $S_{2l}=-\widetilde{K}_{2l+2}(0)/\widetilde{K}_{2l}(0)$ in (2.9), $S_{2l+1}=-\widetilde{K}_{2l+3}^{\prime}(0)/\widetilde{K}_{2l+1}^{\prime}(0)$ in (2.10), one can compute
@@
$=\delta_{m,l}\Lambda_{m+1}+\delta_{m+2,l}\beta_{l}\Lambda_{m+1}+\delta_{m,l+2}\beta_{m}\Lambda_{m-1}+\delta_{m,l}\beta_{m}^{2}\Lambda_{m-1}+\delta_{ml}\alpha_{m}^{2}\Lambda_{m}+\delta_{m+1,l}\alpha_{l}\Lambda_{m+1}+$
$+\delta_{m-1,l}\beta_{m}\alpha_{l}\Lambda_{m-1}+\delta_{m,l+1}\alpha_{m}\Lambda_{m}+\delta_{m,l-1}\alpha_{m}\beta_{l}\Lambda_{m}.$ (2.25)
@@
$=\prod\limits_{j=1}^{m}a_{2j-1}^{2}+a_{2m}^{2}\prod\limits_{j=1}^{m-1}a_{2j-1}^{2}+...+\left(\prod_{k=0}^{\ell}a_{2m-2k}^{2}\right)\left(\prod_{j=1}^{m-\ell-1}a_{2j-1}^{2}\right)+...+\prod_{k=0}^{m-1}a_{2m-2k}^{2}=$
$=\sum\limits_{\ell=-1}^{m-1}\left(\prod_{k=0}^{\ell}a_{2m-2k}^{2}\right)\left(\prod_{j=1}^{m-\ell-1}a_{2j-1}^{2}\right).$ (2.32)
Finally we express the QR transformed coefficients
@@
$\hat{a}^{2}_{2m}=\frac{a^{2}_{2m}a^{2}_{2m+1}}{r^{2}_{2m-1}},\quad\hat{a}^{2}_{2n+1}=r^{2}_{2n+1},$ (2.34)
which ends the proof of Theorem 1.
@@
and the coefficients that are rational functions of $12m+$ 1
@@
$\frac{r_{2m+1}^{2}}{a_{2m+2}^{2}}=\bigg{(}1+\frac{a_{2m+1}^{2}}{a_{2m+2}^{2}}\bigg{)}-\frac{a_{2m+1}^{2}/a_{2m+2}^{2}}{r_{2m-1}^{2}/a_{2m}^{2}}=\bigg{(}1+\frac{a_{2m+1}^{2}}{a_{2m+2}^{2}}\bigg{)}-\frac{a_{2m+1}^{2}/a_{2m+2}^{2}}{\left(1+\dfrac{a_{2m-1}^{2}}{a_{2m}^{2}}\right)-\dfrac{a_{2m-1}^{2}/a_{2m}^{2}}{r_{2m-3}^{2}/a_{2m-2}^{2}}}=$
@@
Then there exist the unique solutions of three term reccurence for $k\geq 1$ (see, for example, [39])
@@
Now we note that in the double-scaling limit $n=HK,m=\frac{\lambda}{2}K,\;,\lambda\in(0,1),\;H\in(0,1/2),\;K\to\infty$ one can find for $k+1\leq m$
@@
One may then employ the WKB method [40] to construct the asymptotic solution of the three-term recurrence relations (2.36) and (2.37).
@@
The first of these equations is nonlinear and of second order. It remains an open question whether this system can be reduced to two coupled nonlinear first-order recurrences, for example by methods analogous to those in [41] and [42].
@@
For illustrative purposes we present first several monic Krawtchouk polynomials $\{\widetilde{K}_{m}(x)\}$ and monic symplectic polynomials $\{G_{m}(x)\}$ for $p=1/2$ in Table 1
$\begin{array}[]{|l|l|}\hline\cr\widetilde{K}_{m}(x)&G_{m}(x)\\ \hline\cr 1&1\\ x&x\\ x^{2}-\frac{K}{4\,n^{2}}&x^{2}-\frac{3\,K+2}{4\,n^{2}}\\ x^{3}-\frac{1}{4}\,x{\left(\frac{3\,K-2}{n^{2}}\right)}&x^{3}-\frac{1}{4}x\frac{\left(15K^{2}-30K+16\right)}{(3K-2)n^{2}}\\ x^{4}-\frac{1}{2}\,x^{2}{\left(\frac{3\,K-4}{n^{2}}\right)}+\frac{3\,K^{2}-6\,K}{16\,n^{4}}&x^{4}-\frac{5}{2}x^{2}\frac{(K-2)}{n^{2}}+\frac{15K^{2}-50K+24}{16n^{4}}\\ x^{5}-\frac{5}{2}\,x^{3}{\left(\frac{K-2}{n^{2}}\right)}+\frac{1}{16}\,x{\left(\frac{15\,K^{2}-50\,K+24}{n^{4}}\right)}&x^{5}-\frac{5}{2}x^{3}\frac{21K^{3}-126K^{2}+224K-96}{\left(15K^{2}-50K+24\right)n^{2}}+\\ &\quad+x\frac{525K^{4}-4200K^{3}+11340K^{2}-11840K+4416}{16\left(15K^{2}-50K+24\right)n^{4}}\\ x^{6}-\frac{5}{4}\,x^{4}{\left(\frac{3\,K-8}{n^{2}}\right)}+\frac{1}{16}\,x^{2}{\left(\frac{45\,K^{2}-210\,K+184}{n^{4}}\right)}-&x^{6}+x^{4}\frac{70-21K}{4n^{2}}+\frac{7}{16}x^{2}\frac{15K^{2}-90K+112}{n^{4}}-\\ \quad-\frac{15\,K^{3}+90\,K^{2}-120\,K}{64\,n^{6}}&\quad-\frac{3}{64}\frac{35K^{3}-280K^{2}+588K-240}{n^{6}}\\ \hline\cr\end{array}$
Table 1. Monic Krawtchouk polynomials and monic symplectic polynomials for $p=1/2$.
3. Asymptotics of the polynomials
In investigating the asymptotic properties of the Christoffel–Darboux kernel (1.13), we are led, via the analysis of equation (2.9), to the necessity of determining the asymptotic behavior of the polynomials $\{\widetilde{K}_{m}\}$.
3.1. Double-scaling asymptotic for Krawtchouk polynomials

### v3

@@ 2.1 Symplectic polynomials
$\sum\limits_{a=0}^{K}\binom{K}{a}p^{a}(1-p)^{K-a}K_{m}(a,p,K)K_{n}(a,p,K)=\binom{K}{n}^{-1}\left(\frac{1-p}{p}\right)^{n}\delta_{nm},$
@@
There is a well known representation of the Krawtchouk polynomials as hypergeometric functions [Koekoek]
@@
The monic version of the $K_{m}(a,p,K)$ is defined as (9.11.4 in [Koekoek])
@@
$P_{m}(a,p,K)$ $=(-1)^{m}m!p^{m}\binom{K}{m}K_{m}(a,p,K)=$
$=(-1)^{m}m!p^{m}\binom{K}{m}\,_{2}F_{1}\left(-m,-a,-K;\frac{1}{p}\right)$ (2.3)
@@
The weight $\widetilde{W}(u^{2})$, defined by (1.8) differs from the weight of Krawthouk polynomials $W$ with $K=2n+2k$ only by a factor $u^{2}$.
Proposition 2 (Christoffel transformation, see [Ismail_2005], sect. 2.7, Theorem 2.7.1).
Monic orthogonal polynomials $G_{m}$ are expressed in terms of the polynomials (2.4) by the formula
@@
[…]
@@
$-\frac{\,{}_{2}F_{1}\left(-m-2,-K/2,-K;\frac{1}{p}\right)}{\,{}_{2}F_{1}\left(-m,-K/2,-K;\frac{1}{p}\right)}\,_{2}F_{1}\left(-m,-K/2-nu,-K;\frac{1}{p}\right)\bigg).$ (2.12)
@@
2.2 QR algorithm and three term relation’s coefficients
Using the results from [BUHMANN1992117], one can derive a three-term recurrence relation for symplectic polynomials. This can be achieved by employing the fact that the Christoffel transformation provides a method for constructing orthogonal polynomials with respect to the measure $x^{2}W(x)dx$ from those that are orthogonal with respect to the positive Borel measure $W(x)dx$.
@@ 2.2 QR algorithm and three term relation’s coefficients
$\int\limits_{-\infty}^{\infty}\pi_{m}(x)\pi_{n}(x)\,d\mu(x)=\delta_{mn},\quad\int\limits_{-\infty}^{\infty}d\mu(x)=1,$ (2.14)
@@
$\int\limits_{-\infty}^{\infty}d\mu(x)=\sum\limits_{i=-K/2}^{K/2}W(i/n)=\sum\limits_{i=-K/2}^{K/2}\binom{K}{K/2+i}p^{K/2+i}(1-p)^{K/2-i}=1.$
@@
$\Omega_{2}(K,p,n)=\int\limits_{-\infty}^{\infty}x^{2}d\mu(x)=\sum\limits_{i=-K/2}^{K/2}(i/n)^{2}\binom{K}{K/2+i}p^{K/2+i}(1-p)^{K/2-i}=$
@@
Now we can determine the coefficients of the three term relation for the polynomials $\{\pi_{m}(x)\}$ from three term relation coefficients $\{\tilde{\alpha}_{m},\tilde{\beta}_{m}\}$ of the monic polynomials $\{\widetilde{K}_{m}\}$ (see [Koekoek])
$x\widetilde{K}_{m}(x)=\widetilde{K}_{m+1}(x)+\tilde{\alpha}_{m}\widetilde{K}_{m}(x)+\tilde{\beta}_{m}\widetilde{K}_{m-1}(x)=\widetilde{K}_{m+1}(x)+$ (2.18)
$+[p(K-m)+m(1-p)-K/2]\widetilde{K}_{m}(x)+\frac{m(K-m+1)}{n^{2}}p(1-p)\widetilde{K}_{m-1}(x).$
@@
$a_{m}=\frac{\kappa_{m-1}}{n\kappa_{m}}=-\frac{\big[m(K-m+1)(p(1-p))\big]^{1/2}}{n}=-\sqrt{\tilde{\beta}_{m}},$ (2.19)
$b_{m}=\left[p(K-m)+m(1-p)-K/2\right]=\tilde{\alpha}_{m}.$ (2.20)
@@
Following the Theorem 3.3 in [BUHMANN1992117], consider the Jacobi matrix of the three term relation coefficient for the orthonormal polynomials with the measure $\mu$. Denote by $\omega$ the natural mapping from the set of all normalized, real, positive Borel measures $\mathcal{B}$ to the set of all singly-infinite, symmetric, tridiagonal irreducible matrices $\mathcal{J}$
@@
that takes every member of $\mathcal{B}$ to the corresponding Jacobi matrix. By the Favard theorem [chihara2011introduction] this mapping is invertible if and only if the underlying Hamburger problem is determinate [akhiezer1965classical]. By the inverse mapping we mean $\omega^{-1}:\mathcal{J}\to\mathcal{B}^{*}$, where $\mathcal{B}^{*}$ is the quotient by the equivalence classes of the measures with identical moment sequences. In our case the measure $d\mu(x)$ defined above has the discrete and finite support, which means that the moment problem for $\text{Im}(\omega)$ is always determinate and the equivalence class of the element $\omega^{-1}(J)$ consists of a single element (see [Simon97], [Christiansen2004IndeterminateMP], [Simon99]). Then the QR algorithm gives the reversible arrows in the diagram
$J^{\widehat{J}}$$d\mu$$d\tilde{\mu}$$\scriptstyle{\tau}$$\scriptstyle{\omega}$$\scriptstyle{\rho}$$\scriptstyle{\omega}$
@@
QRQR factorization of Jacobi matrix $J$ is its decomposition into the product $J=Q_{1}^{T}Q_{2}^{T}...Q_{n-1}^{T}R$, where $Q_{i}$ are block-diagonal antisymmetric matrices, which turn the subdiagonal elements to zero (Givens rotation). For example, letting $r_{0}=\sqrt{a_{1}^{2}+b_{0}^{2}}$, we take $Q_{1}$ to be identity except that the upper left $2\times 2$ block is
@@
The trailing coefficients $b_{i}^{*}$ and $a_{i+1}^{*}$, obtained by applying $i$ successive Givens rotations to $J$, serve as the parameters for the subsequent rotation $Q_{i+1}$. The nontrivial part of $Q_{i+1}$, in its (k+1)(k+1)st and (k+2)(k+2)nd rows and columns, has the form
@@
[…]
Proposition 3.
The squared norms and the coefficients of the three-term recurrence relation for the monic symplectic polynomials are governed by a nonlinear recurrence relation of the form below
$S_{m}L_{m}=\beta_{m}\Lambda_{m-1},$ (2.26)
$L_{m+2}+S_{m}^{2}L_{m}=\Lambda_{m+1}+\beta_{m}^{2}\Lambda_{m-1}+\alpha_{m}^{2}\Lambda_{m},$ (2.27)
$0=\beta_{m}\alpha_{m-1}\Lambda_{m-1}+\alpha_{m}\Lambda_{m}.$ (2.28)
Proof.
From the definition of the Christoffel transformation it follows that
@@
$=\delta_{m,l}\Lambda_{m+1}+\delta_{m+2,l}\beta_{l}\Lambda_{m+1}+\delta_{m,l+2}\beta_{m}\Lambda_{m-1}+\delta_{m,l}\beta_{m}^{2}\Lambda_{m-1}+\delta_{ml}\alpha_{m}^{2}\Lambda_{m}+$
$+\delta_{m+1,l}\alpha_{l}\Lambda_{m+1}+\delta_{m-1,l}\beta_{m}\alpha_{l}\Lambda_{m-1}+\delta_{m,l+1}\alpha_{m}\Lambda_{m}+\delta_{m,l-1}\alpha_{m}\beta_{l}\Lambda_{m}.$ (2.30)
@@
Lemma 1.
The coefficients of three term relation os symplectic polynomials for $p=1/2$ can be computed from (2.19) and (2.20) as
$\hat{b}_{k}=0\;(k\geq 0),$ (2.35)
$\hat{a}^{2}_{2m}=\frac{a^{2}_{2m}a^{2}_{2m+1}}{r^{2}_{2m-1}},\quad\hat{a}^{2}_{2m+1}=r^{2}_{2m+1},$ (2.36)
where
$r_{2m+1}^{2}=a_{2m+2}^{2}+a_{2m+1}^{2}\prod\limits_{j=1}^{m}\frac{a_{2j-1}^{2}}{r_{2j-1}^{2}}.$ (2.37)
Proof.
@@
$=\prod\limits_{j=1}^{m}a_{2j-1}^{2}+a_{2m}^{2}\prod\limits_{j=1}^{m-1}a_{2j-1}^{2}+...+\left(\prod_{k=0}^{\ell}a_{2m-2k}^{2}\right)\left(\prod_{j=1}^{m-\ell-1}a_{2j-1}^{2}\right)+$...+
$+\prod_{k=0}^{m-1}a_{2m-2k}^{2}=\sum\limits_{\ell=-1}^{m-1}\left(\prod_{k=0}^{\ell}a_{2m-2k}^{2}\right)\left(\prod_{j=1}^{m-\ell-1}a_{2j-1}^{2}\right).$ (2.40)
Finally we express the QR transformed coefficients as
@@
$\hat{a}^{2}_{2m}=\frac{a^{2}_{2m}a^{2}_{2m+1}}{r^{2}_{2m-1}},\quad\hat{a}^{2}_{2m+1}=r^{2}_{2m+1},$ (2.42)
which ends the proof of the lemma. ∎
@@
and the coefficients that are rational functions of $2m+1$
@@
$\frac{r_{2m+1}^{2}}{a_{2m+2}^{2}}$ $=\bigg(1+\frac{a_{2m+1}^{2}}{a_{2m+2}^{2}}\bigg)-\frac{a_{2m+1}^{2}/a_{2m+2}^{2}}{r_{2m-1}^{2}/a_{2m}^{2}}=$
$=\bigg(1+\frac{a_{2m+1}^{2}}{a_{2m+2}^{2}}\bigg)-\frac{a_{2m+1}^{2}/a_{2m+2}^{2}}{\left(1+\dfrac{a_{2m-1}^{2}}{a_{2m}^{2}}\right)-\dfrac{a_{2m-1}^{2}/a_{2m}^{2}}{r_{2m-3}^{2}/a_{2m-2}^{2}}}=$
@@
Then there exist the unique solutions of three term reccurence for $k\geq 1$ (see, for example, [deiftorthogonal])
@@
Now we note that in the double-scaling limit $n=HK,\;m=\frac{\lambda}{2}K,\;\lambda\in(0,1),\;H\in(0,1/2),\;K\to\infty$ one can find for $k+1\leq m$
@@
One may then employ the WKB method [Braun1978TMF] to construct the asymptotic solution of the three-term recurrence relations (2.44) and (2.45).
@@
The first of these equations is nonlinear and of second order. It remains an open question whether this system can be reduced to two coupled nonlinear first‐order recurrences, for example by methods analogous to those in [Dzhamay_2018] and [Dzhamay_2020].
@@
Table 1: Monic Krawtchouk polynomials and monic symplectic polynomials for $p=1/2$.
$\widetilde{K}_{m}(x)$ $G_{m}(x)$
11 11
$x$ $x$
$x^{2}-\frac{K}{4\,n^{2}}$ $x^{2}-\frac{3\,K+2}{4\,n^{2}}$
$x^{3}-\frac{1}{4}\,x{\left(\frac{3\,K-2}{n^{2}}\right)}$ $x^{3}-\frac{1}{4}x\frac{\left(15K^{2}-30K+16\right)}{(3K-2)n^{2}}$
$x^{4}-\frac{1}{2}\,x^{2}{\left(\frac{3\,K-4}{n^{2}}\right)}+\frac{3\,K^{2}-6\,K}{16\,n^{4}}$ $x^{4}-\frac{5}{2}x^{2}\frac{(K-2)}{n^{2}}+\frac{15K^{2}-50K+24}{16n^{4}}$
$x^{5}-\frac{5}{2}\,x^{3}{\left(\frac{K-2}{n^{2}}\right)}+\frac{1}{16}\,x{\left(\frac{15\,K^{2}-50\,K+24}{n^{4}}\right)}$ $x^{5}-\frac{5}{2}x^{3}\frac{21K^{3}-126K^{2}+224K-96}{\left(15K^{2}-50K+24\right)n^{2}}+$
   $x\frac{525K^{4}-4200K^{3}+11340K^{2}-11840K+4416}{16\left(15K^{2}-50K+24\right)n^{4}}$
$x^{6}-\frac{5}{4}\,x^{4}{\left(\frac{3\,K-8}{n^{2}}\right)}+\frac{1}{16}\,x^{2}{\left(\frac{45\,K^{2}-210\,K+184}{n^{4}}\right)}$- $x^{6}+x^{4}\frac{70-21K}{4n^{2}}+\frac{7}{16}x^{2}\frac{15K^{2}-90K+112}{n^{4}}$-
   $\frac{15\,K^{3}+90\,K^{2}-120\,K}{64\,n^{6}}$    $\frac{3}{64}\frac{35K^{3}-280K^{2}+588K-240}{n^{6}}$
For illustrative purposes we present first several monic Krawtchouk polynomials $\{\widetilde{K}_{m}(x)\}$ and monic symplectic polynomials $\{G_{m}(x)\}$ for $p=1/2$ in Table  1.
3 Asymptotics of the polynomials
In investigating the asymptotic properties of the Christoffel–Darboux kernel (1.14), we are led, via the analysis of equation (2.9), to the necessity of determining the asymptotic behavior of the polynomials $\{\widetilde{K}_{m}(j/n)\}$.